### v2

# Language Models Do Not Have Human-Like Working Memory

Abstract
While Large Language Models (LLMs) exhibit remarkable reasoning abilities, we demonstrate that they fundamentally lack a core aspect of human cognition: working memory. Human working memory is an active cognitive system that enables not only the temporary storage of information but also its processing and utilization. Without working memory, individuals may produce unreal conversations, exhibit self-contradiction, and struggle with tasks requiring mental reasoning. Existing evaluations using N-back or context-dependent tasks fail as they allow LLMs to exploit accessible context rather than retain latent information. We introduce three novel tasks—(1) Number Guessing, (2) Yes-No Deduction, and Math Magic—that isolate internal representation from external context. Across seventeen frontier models spanning four major model families, we consistently observe irrational or contradictory behaviors, highlighting LLMs’ inability to retain and manipulate latent information. Our work establishes a new benchmark for evaluating working memory in LLMs and identifies this deficit as a critical obstacle to artificial general intelligence. Code and prompts for the experiments are available at https://github.com/penguinnnnn/LLM-Working-Memory.
Figure 1: When LLMs say they already have a number in mind, and it is not 4, how can we know whether LLMs are lying, or even thinking of nothing?
1 Introduction
Large Language Models (LLMs) have demonstrated strong performance across diverse domains, including legal advice Guha et al. (2023), education Wen et al. (2024), translation Jiao et al. (2023), and healthcare Yang et al. (2024b). Beyond task performance, recent studies have examined their intrinsic behaviors, often drawing parallels with human cognition, such as personality Huang et al. (2024b), emotion Huang et al. (2024a), empathy Sorin et al. (2024), theory-of-mind Liu et al. (2024b), and values Wang et al. (2024b).
One such behavior is memory, which has attracted increasing attention from both industry and research communities. OpenAI was the first to introduce a memory module in ChatGPT (February 2024). [^1] https://openai.com/index/memory-and-new-controls-for-chatgpt/ By mid-2025, xAI, Anthropic, and Google integrated memory into Grok, [^2] https://x.com/grok/status/1912670182012801156 Claude, [^3] https://www.anthropic.com/news/claude-4 and Gemini, [^4] https://blog.google/products/gemini/temporary-chats-privacy-controls/ respectively. Sam Altman even emphasized that “people want memory,” identifying it as a central feature of the forthcoming GPT-6.555https://futurism.com/disastrous-gpt-5-sam-altman-hyping-up-gpt-6 On the research side, memory modules are also actively studied Zhang et al. (2025); Hatalis et al. (2023); Xu et al. (2025).
What is memory? Atkinson & Shiffrin (1968) categorized memory by retention timescale: sensory (1ms–2s), short-term (seconds), and long-term (hours to lifetime). Long-term memory enables information storage over extended periods, whereas short-term, or Working Memory, maintains and manipulates information during complex tasks such as reasoning, comprehension, and learning Baddeley & Hitch (1974). These distinctions have also been adopted in machine learning, corresponding to representational learning for raw inputs (sensory memory), in-context computation at test time (working memory), and access to external databases (long-term memory) Weng (2023). As a central aspect of human cognition, memory has been widely benchmarked Du et al. (2025). However, working memory remains relatively underexplored.
Existing studies often adopt the N-back task Kirchner (1958) to assess LLMs’ working memory Gong et al. (2024); Zhang et al. (2024). Yet, a fundamental limitation arises: the critical information for correct responses remains accessible in the model’s input context. Unlike humans, who cannot revisit prior steps, LLMs can simply attend to earlier tokens within their context window. Fig. 2 illustrates this discrepancy. To more faithfully evaluate working memory, it is necessary to design experiments where the key information is not explicitly present in the context. Imagine the following scenario: You select a number between one and ten. When ready, you are asked, “Is the number greater than five?” If you answer, observers can reasonably infer that the number has entered your conscious awareness (i.e., your working memory), since clear perception is necessary to perform the comparison and provide a response. [^6] A possibility remains that the response was given by chance.
We ask: Do LLMs possess human-like working memory beyond evaluations that merely test their ability to attend to visible information? AI models are often viewed as reasoning in two modes: (1) the token space over sequences Wei et al. (2022); Yao et al. (2023), and (2) the latent space over activations Hao et al. (2024); Geiping et al. (2025). Our focus is on the second mode, where evaluating working memory provides insight into whether models can hold and manipulate latent concepts without explicit externalization. Success in this capacity could enhance reasoning without reliance on Chain-of-Thought (CoT), as it directly tests the model’s ability to maintain objects and concepts internally. Conversely, deficits in working memory impair information processing in humans Gruszka & Nkecka (2017); Cowan (2014), and in LLMs manifest as unrealistic outputs, self-contradictions, and failures on tasks requiring mental manipulation.
The central challenge in designing such evaluations is epistemic: how can we demonstrate the absence of internal memory when we cannot directly observe a model’s mind? To address this limitation, we design three experiments—(1) Number Guessing, (2) Yes–No Deduction, and (3) Math Magic—that test whether LLMs can internally maintain information that is not explicitly present in the input context. Our experiments span 17 state-of-the-art LLMs, both proprietary and open-source, and reveal a consistent inability to perform such tasks. The failure is uniform across model families (GPT Hurst et al. (2024), Qwen Yang et al. (2024a), DeepSeek Liu et al. (2024a); Guo et al. (2025), LLaMA Grattafiori et al. (2024)) and reasoning approaches (CoT Wei et al. (2022) or o1-like reasoning Jaech et al. (2024)).
2 Background
Figure 2: a. An illustration of how “N-Back” tasks are performed. b. Humans see the stimuli one after one, forcing them to put the information in working memory. c. Researchers put all stimuli into context, enabling LLMs to easily find the answers.
2.1 How Working Memory is Evaluated
Human working memory is typically assessed using behavioral paradigms that require individuals to maintain and update information over short intervals. Common examples include the digit span task Miller (1956), where participants recall sequences of numbers of increasing length, and the N-back task Kirchner (1958), which requires identifying whether the current stimulus matches one presented $N$ steps earlier. These tasks are widely used because they probe the ability to maintain and manipulate information that is no longer externally visible, thereby capturing the essence of working memory function.
In the context of LLMs, working memory has been widely discussed. However, most studies use the term to describe an LLM’s capacity to process information within a fixed context window Li et al. (2023); Guo et al. (2023). Gong et al. (2024); Zhang et al. (2024) evaluate LLMs using N-back tasks. As noted in §1, such human-designed tests are not directly valid for LLMs, since models can simply attend to retained tokens in their context window rather than actively maintaining information.
2.2 Existing Engineering Solutions
A separate line of work introduces explicit engineering solutions by equipping LLMs with external memory modules Hu et al. (2024); Zeng et al. (2024). For example, Wang et al. (2024a) incorporate symbolic working memory to enhance reasoning, while Kang et al. (2024) use it to improve training efficiency. Other approaches implement scratch spaces Lanchantin et al. (2023) or internal CoT mechanisms Jaech et al. (2024), which may partially mitigate working memory limitations identified in this paper. However, these methods do not address whether LLMs possess an intrinsic working memory capacity, analogous to humans’ ability to mentally simulate and manipulate objects. The distinction can be illustrated with an analogy: modern calculators make arithmetic trivial, yet schools continue to assess addition, subtraction, multiplication, and division. The purpose is not solely to solve problems efficiently, but to reveal individuals’ underlying cognitive abilities. Similarly, while engineering techniques can extend an LLM’s effective memory, we are ultimately interested in whether the model itself has the basic cognitive capability of working memory. Without it, a model may function adequately through external tools, but its intrinsic reasoning ability remains fundamentally limited.
3 Number Guessing
(a) Probabilities for OpenAI GPT models.
(b) Probabilities for OpenAI o-series models.
(c) Probabilities for Meta LLaMA models.
(d) Probabilities for Qwen and DeepSeek models.
Figure 3: Probabilities of model answering “Yes” for each number from one to ten.
Consider a number-guessing game in which a human participant privately selects a number between one and ten, and we ask whether the chosen number is one. By repeating this process hundreds of times, we can estimate the probability of the participant selecting the number one. Repeating the procedure for each possible number yields an empirical distribution over the participant’s choices. It is worth noting that, if the participant actually selects a number and responds honestly, the estimated probabilities across all numbers should sum to one.
Figure 4: Probabilities of GPT-4o-2024-08-06 answering “Yes” for each number in different ranges.
Setup.
Leveraging this property, we design an experiment to probe LLMs. In each trial, the model is given a fixed prompt: “User: Think of an integer between 1 and 10, but don’t say it to me. Assistant: Got it! I’ve thought of an integer between 1 and 10. What’s next?” The model is then prompted 200 times for each integer from 1 to 10 with queries such as “Is the number you’re thinking of 7? Answer Yes or No.” We record the frequency of “Yes” responses for each number and treat these frequencies as probabilities. If the sum of these probabilities deviates significantly from one, it suggests that LLMs either do not think of a specific number or lie to users. All models are configured with a temperature of 1.0 and a top‑p value of 1.0.
Table 1: The sum of probabilities of each model responding “Yes” for all numbers from one to ten. Color intensity reflects proximity to one: red indicates values closer to zero, while blue signifies values greater than one.
Model Sum
GPT-4o-Mini-2024-07-18 0
GPT-4o-2024-05-13 0
GPT-4o-2024-08-06 1.085
GPT-4o-2024-11-20 0
GPT-4.1-2025-04-14 0
o1-Mini-2024-09-12 0.005
o3-Mini-2025-01-31 0.205
o4-Mini-2025-04-16 0.030
LLaMA-3.3-70B-Instruct-Turbo 0.045
LLaMA-3.1-8B-Instruct-Turbo 0.980
LLaMA-3.1-70B-Instruct-Turbo 0.465
LLaMA-3.1-405B-Instruct-Turbo 1.195
Qwen2.5-7B-Instruct-Turbo 0
Qwen2.5-72B-Instruct-Turbo 0
QwQ-32B 0.005
DeepSeek-V3 0
DeepSeek-R1 0.640
Findings.
Fig. 3 presents the probabilities of “Yes” responses in each model for numbers from one to ten. Two key observations emerge: (1) Most LLMs never produce a “Yes” response; “No” dominates across models. This suggests that LLMs generate unrealistic distributions, implying an inability to produce realistic responses. Given that LLMs generally follow instructions and do not deliberately deceive, we attribute this behavior to their failure to internally “think of” a number. (2) When LLMs do respond affirmatively, they exhibit a marked preference for the number seven. This tendency mirrors patterns observed in humans Miller (1956); Kubovy & Psotka (1976).
We quantify LLM performance on this task using the sum of probabilities. A value closer to one indicates better model performs. Table 1 reports these sums for each model. Several observations emerge: (1) Newer models do not necessarily outperform older ones. Within the GPT family, the 0806 version of GPT-4o (the current target of the “gpt-4o” API) achieves the best performance, surpassing both the 1120 version and GPT-4.1. Similarly, LLaMA-3.3 underperforms relative to LLaMA-3.1. (2) Using CoT reasoning does not improve performance. Models employing such strategies—o1, o3, o4, QwQ, and DeepSeek-R1—fail to produce probability sums closer to one. (3) Overall, LLaMA-3.1 performs best, with the 8B variant outperforming both the 405B and 70B versions.
Table 2: The sum of probabilities of GPT-4o-2024-08-06 responding “Yes” for all numbers in different ranges.
Number Sum
2 0
3 2.180
5 1.600
7 0.920
9 2.025
10 1.085
20 0.080
30 0
40 0.005
Other Number Ranges.
We extend our experiments to include a broader range of numbers. Given that GPT-4o-2024-08-06 performs best among the OpenAI models, we focus on its behavior across different numerical ranges. Table 2 reports the summed probabilities for each range, while Fig. 4 illustrates the probability of individual numbers. Our findings reveal two key patterns: (1) For smaller ranges such as 3, 5, and 9, the model exhibits a strong bias toward answering “Yes,” with the probability sum significantly exceeding one. In contrast, for larger ranges like 20, 30, and 40, “Yes” responses are rare. (2) When the model does produce a “Yes” response, it frequently corresponds to numbers ending in seven (e.g., 7, 17, 37), as shown in Fig. 4.
In conclusion, LLMs fail to generate correct distributions, suggesting they struggle to produce realistic responses. This limitation likely arises from their inability to internally think of numerical values without explicit contextual cues.
4 Yes-No Deduction
Table 3: Objects ordered by the five properties (smallest to largest).
Volume Length Weight Density
Coffee bean Rice Coin Air
Dice Paperclip Spoon Wood
Golf ball Credit card Watch Ice
Soda can Pencil Smartphone Water
Soccer ball Laptop Bottle of water Plastic
Microwave oven Baseball bat Dictionary Glass
Washing machine Guitar Cat Iron
Bathtub Door Bicycle Copper
Car Apple tree Television Silver
School bus Coconut tree Refrigerator Gold
Shipping container Tennis court Tiger Hardness
Olympic swimming pool Swimming pool Cow Marshmallow
Boeing 747 Football field Rhino Rubber eraser
Titanic Skyscraper Elephant Brick
Great Pyramid of Giza Mount Everest Train Hammer
Diamond ring
“Yes-No” is a social deduction game commonly used to train human reasoning, classification, and questioning skills. In this game, one player privately selects an object, while the opponent asks yes–no questions (e.g., “Is the object heavier than an elephant?”) to progressively narrow the possibilities and ultimately guess the object. Consider the decision-making process of the player answering questions. Upon receiving a question, the player simply compares the imagined object with the queried attribute and responds accordingly. Note that, humans do not typically recall all previous questions and answers to ensure consistency; instead, they answer each new question based on immediate judgment, without explicitly checking self-contradiction with prior responses.
We hypothesize that if LLMs lack the working memory to temporarily retain an imagined object, they can only respond to questions by checking consistency with their prior answers. As the number of questions increases, maintaining consistency becomes increasingly difficult. Consequently, performance on this task is closely tied to the model’s ability for long-context reasoning. In this section, we present an experiment in which the LLM is first instructed to imagine an object, followed by a sequence of comparative questions involving that object and other reference objects. The goal is to assess whether the model produces self-contradictions when answering sequential questions about the imagined object. For instance, the model might initially answer “Yes” to “Is the object heavier than an elephant?” but later also respond “Yes” to “Is the object lighter than a cat?”, thereby contradicting itself.
Setup.
We predefined five sets of objects that are commonly regarded as comparable with respect to five properties: volume, length, weight, density, and hardness. In total, 60 distinct objects are included, as listed in Table 3, ordered by the corresponding property. For each question, one property is randomly selected, followed by an object from the corresponding object list. The LLM is then prompted to assess whether the object it has in mind is comparative than the given object, where the comparative form can vary in direction (e.g., bigger or smaller for volume). In each trial, the model is continuously presented with up to 250 such questions. We record the number of questions completed before the model exhibits a self-contradiction. If no contradiction is observed across all 250 questions, the trial is considered a Pass. Each model is tested with 200 trials.
Table 4: Count of failures of GPT-4o-Mini-2024-07-18 and GPT-4o-2024-08-06.
Model Failure V W L D H
GPT-4o-Mini 200 12 46 49 52 41
GPT-4o 173 21 42 57 27 26
Findings.
Table 4 presents the number of failed trials for the two models. The smaller model (GPT-4o-Mini) consistently fails, indicating that it always produces self-contradiction. As model capacity increases, GPT-4o successfully passes 27 out of 200 trials. This result supports our hypothesis that model performance on this task depends on the their long-context processing ability—an outcome that would be unlikely if the models possessed the working memory for maintaining imagined objects.
Figure 5 presents histograms of the number of questions each model completes before exhibiting self-contradiction. The distribution for GPT-4o-Mini peaks in the 20–30 range, whereas GPT-4o peaks in the 30–40 range. Moreover, GPT-4o demonstrates a higher frequency of completions in the 80–130 range compared to GPT-4o-Mini. Notably, the types of properties that lead to self-contradictions differ between the two models: GPT-4o-Mini more frequently fails on density and hardness, while GPT-4o shows greater robustness on these attributes.
Table 5: Ablation studies using GPT-4o-2024-08-06.
Model Failure V W L D H
Hint 194 37 39 60 37 21
All 158 18 29 21 55 35
Hint + All 145 13 32 34 46 20
Ablation.
To ensure that the observed failures are not simply due to LLMs’ inability to rank objects by the five properties, we conduct the following ablation studies: (1) Hint: At the beginning of the prompt, we provide GPT-4o with the object rankings defined in Table 3. (2) All: For each question, we specify the target object $\mathcal{O}$ by stating explicitly that “the object you are considering is $\mathcal{O}$.” (3) Hint + All: We combine the above two settings. Results are shown in Table 5. Two key findings emerge: (1) Even when given hints that guarantee knowledge of the correct ranking, GPT-4o still frequently fails, suggesting that the task critically depends on long-context reasoning. (2) When the object identity is provided explicitly, error rates decrease, likely because the long-context reasoning task is reduced to a short-context one.
In conclusion, LLMs exhibit self-contradictions (e.g., claiming an object is both larger than a car and smaller than a soccer ball) as the number of queries increases. This behavior suggests their reliance on long-context reasoning rather than a proper working memory.
(a) GPT-4o-Mini-2024-07-18.
(b) GPT-4o-2024-08-06.
Figure 5: The histogram of the number of questions where the two models show self-contradiction.
5 Math Magic
Let’s play a game before presenting the final part of our experiments. Think of a three-digit number in which the hundreds and units digits differ (e.g., abc). Reverse the digits to form a new number (e.g., abc $\rightarrow$ cba), and subtract the smaller of the two numbers from the larger. Then, reverse the result—if it is a two-digit number, prepend a zero (e.g., 67 $\rightarrow$ 067 $\rightarrow$ 760). Add this reversed number to the previous result What if we can tell that your result is 1089? It is math who tells us the secret. Such magic, often referred to as “math magic” or “math mentalism,” rely on carefully constructed transformations that yield predictable outcomes, allowing magicians to perform mind-reading.
To perform this task, humans must internally generate or retain numerical information (e.g., remember a poker card) and manipulate it mentally without external cues. This process relies on working memory, a cognitive function that LLMs may lack. Without this capacity, LLMs are prone to errors in mathematical tricks, failing to produce the intended outcomes. Accordingly, we can design complex math magics to evaluate whether LLMs possess working memory–like capabilities.
Table 6: Operations in the math magic in our experiment. the random number $a$ ranges from 1 to 7, while the random numbers $b$ and $c$ range from 1 to 3.
Role Content
User Think of 4 integers between 1 and {NUMBER} in order, but don’t tell me.
Assistant Okay! I’ve got 4 numbers. What’s next?
User In order, append the same 4 numbers after the original ones.
Assistant Understood! Now I have 8 numbers. What’s next?
User Move the first {random_number_a} numbers to the end.
Assistant Got it! Now I have moved the numbers. What’s next?
User Take the first 3 numbers and insert them anywhere in the middle.
Assistant Okay! The first 3 numbers are placed somewhere in the middle. What’s next?
User Set the first number aside. We don’t need it for now.
Assistant Understood! Now I have 7 numbers. What’s next?
User Take the first {random_number_b} numbers and insert them anywhere in the middle.
Assistant Got it. The first {random_number_b} numbers are placed somewhere in the middle. What’s next?
User Remove the first {random_number_c} numbers. We will never need it anymore.
Assistant Okay! Now I have {7 - random_number_c} numbers. What’s next?
User Move the first number to the end. Repeat this seven times.
Assistant Understood! Now my sequence has rearranged. What’s next?
User Remove the second number, and then move the first number to the end. Repeat this {6 - random_number_c} times.
Assistant Got it! Now I have only 1 number. What’s next?
User Tell me what the last remaining number is. Do you remember the number you set aside at the beginning? Tell me what that number was.
Setup.
Since the 1089 game is well known and likely included in the training data of LLMs, our preliminary experiments show that LLMs can recognize the game and thus invalidate the evaluation. Therefore, to more effectively assess LLMs’ multi-step mental manipulation, we select a more complicated math magicbased on the Josephus Problem Schumer (2002). In this task, participants are asked to imagine four numbers and perform a sequence of operations, including duplication, rotation, and removal. The procedure is shown in Table 6. Ultimately, only two numbers remain, and mathematical constraints guarantee they are identical. In our experiment, LLMs are prompted to secretly think of four numbers and mentally execute the sequence of randomized operations. We measure the proportion of trials (out of 150) in which the LLM completes the task correctly.
Table 7: LLM performance on the math magic task.
Model Count Acc (%)
GPT-4o-Mini-2024-07-18 0/150 0.0
GPT-4o-2024-05-13 4/150 2.7
GPT-4o-2024-08-06 3/150 2.0
GPT-4o-2024-11-20 0/150 0.0
GPT-4.1-2025-04-14 - -
LLaMA-3.3-70B-Instruct-Turbo 7/150 5.7
LLaMA-3.1-8B-Instruct-Turbo 20/150 13.3
LLaMA-3.1-70B-Instruct-Turbo 7/150 5.7
LLaMA-3.1-405B-Instruct-Turbo 39/150 26.0
Qwen2.5-7B-Instruct-Turbo 8/150 5.3
Qwen2.5-72B-Instruct-Turbo 2/150 1.3
DeepSeek-V3 4/150 2.7
(a) LLMs without CoT. GPT-4.1-2025-04-14 fails to complete most of the cases, incorrectly assuming that the necessary numerical inputs are missing.
Model w/ CoT or LRM Count Acc (%)
GPT-4o-Mini-2024-07-18 5/150 3.3
GPT-4o-2024-05-13 26/150 17.3
GPT-4o-2024-08-06 31/150 20.7
GPT-4o-2024-11-20 - -
LLaMA-3.3-70B-Instruct-Turbo 25/150 16.7
Qwen2.5-7B-Instruct-Turbo 49/150 32.7
Qwen2.5-72B-Instruct-Turbo 37/150 24.7
DeepSeek-V3 48/150 32.0
o1-Mini-2024-09-12 75/150 50.0
o3-Mini-2025-01-31 145/150 96.7
o4-Mini-2025-04-16 54/150 36.0
QwQ-32B 135/150 90.0
DeepSeek-R1 150/150 100
(b) LLMs with CoT and Large Reasoning Models. GPT-4o-2024-11-20 consistently fails this task.
Findings.
Table 7a reports the accuracy of prompting models to output the two numbers directly. Most LLMs perform poorly on this task, with notably higher accuracy observed in the LLaMA model family. This finding aligns with results from the number guessing game shown in Table 1, where LLaMA models generate more realistic distributions than other models. Together, these consistent results suggest that current LLMs generally lack the working memory required for such tasks.
We further examine whether CoT prompting improves performance on this task. Table 7b presents the results of prompting models to reason step by step, as well as the performance of o1-like long reasoning models (LRMs). Base models prompted to reason step-by-step achieve 10–30% accuracy—substantially higher than without CoT. DeepSeek-R1 attains 100% accuracy, and other LRMs also perform well, indicating that extended token-level reasoning substantially improves accuracy. However, consistent with our number-guessing experiment, models exhibit a strong preference for the number seven, which appears to influence performance—especially within the o-series. Among o1-Mini’s 75 correct predictions, 50 (66.7%) select 7; for o3-Mini and o4-Mini, the counts are 68/145 (46.9%) and 37/54 (68.5%), respectively. These findings suggest that models may not fully solve the task but instead exploit biased number preferences: notably, o3-Mini—being least likely to guess 7—outperforms the other two o-series models. In conclusion, current LLMs struggle with a class of tasks demanding internal object imagination and mental manipulation, relying on an effective working memory.
6 Conclusion
In this study, we present three carefully designed experiments to investigate whether LLMs possess human-like working memory for processing information and generating responses. Across all experiments, the results reveal a consistent pattern: LLMs do not exhibit behavior indicative of a functional working memory. They fail to internally represent and manipulate transient information across multiple reasoning steps, relying instead on the immediate prompt context. Even advanced prompting strategies, such as CoT prompting, yield only marginal improvements on tasks requiring internal state management. This limitation leads to unrealistic responses, self-contradictions, and an inability to perform mental manipulations. We hope these findings encourage future research on integrating working memory mechanisms into LLMs to enhance their reasoning capabilities and human-likeness.
The Use of Large Language Models
LLMs were employed in a limited capacity for writing optimization. Specifically, the authors provided their own draft text to the LLM, which in turn suggested improvements such as corrections of grammatical errors, clearer phrasing, and removal of non-academic expressions. LLMs were also used to inspire possible titles for the paper. While the system provided suggestions, the final title was decided and refined by the authors and is not directly taken from any single LLM output. In addition, LLMs were used as coding assistants during the implementation phase. They provided code completion and debugging suggestions, but all final implementations, experimental design, and validation were carried out and verified by the authors. Importantly, LLMs were NOT used for generating research ideas, designing experiments, or searching and reviewing related work. All conceptual contributions and experimental designs were fully conceived and executed by the authors.